# Frenet Immersed Finite Element Spaces on Triangular Meshes

### v1

@@ 4.2. Efficient generation of matrices and vectors
$S_{K,q}$ $=$ $\sum_{s=\pm}\beta^{s}V_{\eta}^{s}W^{s}(V_{\eta}^{s})^{T}+\beta^{s}\rho^{2}V_{\xi}^{s}W^{s}(V_{\xi}^{s})^{T}$ (4.45)
$=$ $\sum_{s=\pm}\beta^{s}(C^{s})^{T}L_{\eta}^{s}W^{s}(L_{\eta}^{s})^{T}C^{s}+\beta^{s}\rho^{2}(C^{s})^{T}L_{\xi}^{s}W^{s}(L_{\xi}^{s})^{T}C^{s}.$
Once again, the generation of the stiffness matrix requires the evaluation of the first derivative of polynomial basis $R_{i}$ at the quadrature points, there is no need to evaluate the piecewisely defined IFE basis functions.
The construction of the local source vector $\mathbf{f}$ on an interface element follows a similar procedure. In fact,
${\bf f}\approx{\bf f}_{q}=\sum_{s=\pm}V^{s}W^{s}\mathbf{r}^{s}=\sum_{s=\pm}(C^{s})^{T}L^{s}W^{s}\mathbf{r}^{s},\qquad\text{where}\quad\mathbf{r}^{s}=\Big(f(\mathbf{x}_{1}^{s}),f(\mathbf{x}_{2}^{s}),\cdots,f(\mathbf{x}_{n_{q}^{s}}^{s})\Big)^{T}.$ (4.46)
@@ 5. Reconstruction of Frenet-IFE Basis on Triangles
The GC-IFE bases in the general construction (4.32) focus primarily on enforcing the jump conditions, but not on computational performance when used to approximate interface problems. In this section, we show that the reconstruction procedure proposed in [2025AdjeridLinMeghaichi2] is also applicable to constructing a basis for the GC-IFE space on a triangular mesh with improved computational performance; in particular, the associated mass matrix is optimally conditioned.
For each interface element $K$, we write Frenet-IFE basis functions $\hat{\lambda}_{j}$ produced by the general construction procedure in the previous as follows
@@
We include $C_{j}^{-},C_{j}^{+}$ in this notation to emphasize that these IFE basis are completely determined by $C_{j}^{-},~C_{j}^{+}$ once the polynomial basis $R_{j}$ for $\mathbb{P}_{m}$ is chosen. Also, we denote the basis of $\hat{\mathcal{V}}^{m}_{\hat{\beta}}(\hat{K}_{F})$ by
$\hat{B}(C)=\hat{B}(C^{-},C^{+}):=\Big\{\hat{\lambda}_{j}(\eta,\xi;C_{j}^{-},C_{j}^{+}):j=1,\cdots,d_{m}\Big\}.$
Following [2025AdjeridLinMeghaichi2], we proceed to produce a reconstructed Frenet–IFE basis such that the resulting mass matrix has optimal conditioning. We first recall the Lemma 1 from [2025AdjeridLinMeghaichi]:
Lemma 5.1.
@@
Let $\tilde{C}=CQ$, by (4.42), the mass matrix $)M_{K}(\hat{B}(\tilde{C}$)) on interface element $K$ associated with the basis $\hat{B}(\tilde{C})=\hat{B}(\tilde{C}^{-},\tilde{C}^{+})=\hat{B}(C^{-}Q,C^{+}Q)$ can be approximated by
$M_{K,q}(\hat{B}(\tilde{C}))=\sum_{s=\pm}(\tilde{C}^{s})^{T}L^{s}W^{s}(L^{s})^{T}\tilde{C}^{s}=\sum_{s=\pm}(C^{s}Q)^{T}L^{s}W^{s}(L^{s})^{T}C^{s}Q=Q^{T}M_{K,q}(\hat{B}(C))Q.$ (5.47)
@@
We choose $Q$ such that the mass matrix becomes the identity matrix. Let $V_{1}\Lambda V_{1}^{T}=M_{K,q}(\hat{B}(C))$ be the singular value decomposition of the matrix $M_{K,q}(\hat{B}(C))$. Let $Q_{1}=V_{1}\Lambda^{-1/2}$ and $C^{(1)}=CQ_{1}$. Then we have
$M_{K,q}(\hat{B}(C^{(1)}))=M_{K,q}(\hat{B}(CQ_{1}))=Q_{1}^{T}M_{K,q}(\hat{B}(C))Q_{1}=\Lambda^{-1/2}V_{1}^{T}M_{K,q}(\hat{B}(C))V_{1}\Lambda^{-1/2}=I.$ (5.48)
@@
The resulting approximate mass matrix $M_{K,q}(\hat{{B}}(C^{(1)}))$ achieves an optimal condition number in theory. However, in practical finite-precision computations, this optimal conditioning may not be preserved, especially for GC-IFE bases constructed using higher-degree polynomials, as corroborated by the numerical examples in the next section.
@@
This approach uses a factorization of matrix $M_{K,q}(\hat{B}(C))$ as follows. Denote $n_{q}=n_{q}^{-}+n_{q}^{+}$, then
@@
then, by (4.42) we can express $M_{K,q}(\hat{B}(C))$ as
$M_{K,q}(\hat{B}(C))=V(\hat{B}(C))^{T}WV(\hat{B}(C)),$ (5.50)
@@
Let $\tilde{V}(\hat{B}(C))=W^{1/2}\,V(\hat{B}(C))$, then we can factorize the matrix $M_{K,q}(\hat{B}(C))$ as follow
$M_{K,q}(\hat{B}(C))=\tilde{V}(\hat{B}(C))^{T}\tilde{V}(\hat{B}(C)).$
@@
$M_{K,q}(\hat{{B}}(C^{-}Q_{2},\,C^{+}Q_{2}))=Q_{2}^{T}M_{K,q}(\hat{{B}}(C))Q_{2}=\Sigma^{-1}V_{2}^{T}\tilde{V}^{T}\tilde{V}V_{2}\Sigma^{-1}=I_{d_{m}\times d_{m}}.$
Both reconstruction approaches make approximate mass matrix optimally conditioned. Moreover, when the eigenvalues of $M_{K,q}(\hat{{B}}(C))$ are distinct and computations are exact, the two approaches are mathematically equivalent. In this case, they generate the same GC-IFE basis, i.e.,
@@
However, as already emphasized in [2025AdjeridLinMeghaichi2], in finite-precision computations, the two reconstruction approaches may behave differently. This is because the condition number of $M_{K,q}(\hat{B}(C))$ is the square of the condition number of $\tilde{V}(\hat{B}(C))$ since $\Lambda=\Sigma^{2}$. As a result, $M_{K,q}(\hat{B}(C))$ is typically more ill-conditioned than $\tilde{V}$, and the numerical accuracy of the SVD in Reconstruction Approach 1 (and therefore of $Q_{1}$) is limited by round-off errors in the small singular values of $M_{K,q}(\hat{B}(C))$. Consequently, the matrix $Q_{1}$ obtained numerically may deviate substantially from its exact counterpart. In contrast, the SVD of $\tilde{V}(\hat{B}(C))$ in Reconstruction Approach 2 is less sensitive to round-off errors because $\tilde{V}$ has a smaller condition number. Thus the transformation matrix $Q_{2}$ is generally more accurate.
@@ 6. Numerical Experiments
In this section, we present a set of numerical experiments to demonstrate the effectiveness of the proposed geometry-conforming immersed finite element spaces on triangular meshes.
In these numerical examples, the computational domain is $\Omega=(-1,1)^{2}$ with a circular interface $\Gamma=\{|\mathbf{x}|=r_{0}\}$ where $r_{0}=1/\sqrt{3}$, which partitions $\Omega$ into two subdomains $\Omega^{-}=\{|\mathbf{x}|<r_{0}\}$ and $\Omega^{+}=\{|\mathbf{x}|>r_{0}\}$. The diffusion coefficient $\beta(\mathbf{x})$ is piecewise constant $\beta(\mathbf{x})|_{\Omega^{\pm}}=\beta^{\pm}$ where $\beta^{-}=1$ and $\beta^{+}$ is varied in the tests. The exact solution of the boundary problem (1.1) and (1.2) is :
$u(\mathbf{x})=\begin{cases}\dfrac{1}{\beta^{-}}\cos(\pi|\mathbf{x}|^{2}),&\mathbf{x}\in\Omega^{-},\\[5.0pt] \dfrac{1}{\beta^{+}}\cos(\pi|\mathbf{x}|^{2})+\cos(\pi r_{0}^{2})\left(\dfrac{1}{\beta^{-}}-\dfrac{1}{\beta^{+}}\right),&\mathbf{x}\in\Omega^{+},\end{cases}$
which satisfies the interface jump conditions (1.3)-(1.4) as well as the Laplacian extended jump condition (1.5). We examine three types of bases: $\{\hat{\lambda}_{i}\}_{i=1}^{d_{m}}$ produced by the general construction procedure, $\{\hat{\lambda}_{i}^{(1)}\}_{i=1}^{d_{m}}$ by Reconstruction Approach 1, and $\{\hat{\lambda}_{i}^{(2)}\}_{i=1}^{d_{m}}$ by Reconstruction Approach 2. Polynomial degrees up to $p=9$ are tested.
@@
Figure 3. A structured triangular mesh and an unstructured mesh
6.1. Conditioning of Mass Matrices
We first investigate the conditioning of the mass matrices associated with different choices of basis functions. Local mass matrices are computed on all interface elements of an unstructured triangular mesh with $h=\max\{h_{K}:K\in\mathcal{T}_{h}\}=1/10$, and the maximum condition number over all interface elements is reported in Table 1. Three types of basis functions are considered. For the general construction, the condition numbers grow rapidly as both the polynomial degree and the jump ratio increase. In contrast, both reconstruction approaches appear largely insensitive to these variations. Moreover, the second reconstruction approach demonstrates greater robustness for higher-degree polynomials. When the degree exceeds 7, the condition numbers associated with the first reconstruction increase noticeably, whereas those of the second reconstruction remain well controlled.
Table 1. Largest condition number of local mass matrices among all interface elements on an unstructured triangular mesh with $h=1/10$.
@@ 6.1. Conditioning of Mass Matrices
1 4.75E+01 1 1 4.55E+03 1 1 4.45E+05 1 1
2 2.75E+03 1 1 4.15E+04 1 1 4.16E+06 1 1
3 2.70E+05 1 1 1.09E+06 1 1 1.07E+08 1 1
4 9.96E+06 1 1 4.33E+07 1 1 2.01E+09 1 1
5 6.32E+08 1 1 1.56E+10 1 1 1.48E+11 1 1
6 4.13E+10 1 1 6.24E+11 1 1 2.29E+13 1 1
7 1.66E+12 1 1 3.37E+13 1 1 1.14E+15 1 1
8 1.15E+14 1 1 3.59E+15 1.59 1 1.43E+17 4.69E+01 1
9 2.57E+17 1.20E+02 1 9.75E+18 1.82E+03 1 3.77E+20 2.00E+05 1
Table 2 reports the conditioning of the global mass matrices. Since the global mass matrix is block diagonal, its condition number is equal to the maximum condition number among all elements in the mesh. For each polynomial degree $1\leq m\leq 9$, we report the condition numbers on three different meshes. When $h=1/10$, the condition number obtained from the general construction coincides with that in Table 1, because the largest condition number occurs on the interface elements. Moreover, no apparent growth trend is observed as the mesh is refined. This indicates that the geometric configuration of the interface elements has a much stronger impact on the conditioning than the element size itself.
In contrast, both reconstruction approaches yield identical condition numbers across all meshes for polynomial degrees up to $m=8$. In these cases, the observed condition numbers are entirely determined by the non-interface elements, where a finite element basis with a certain orthogonal property [2008HesthavenWarburton] is employed, since the condition numbers on all interface elements are equal to 1, as reported in Table 1. For $m=9$ with $\beta^{+}=1000$, the condition number of the first reconstruction is dominated by the interface element, whereas the second reconstruction remains robust.
Table 2. Condition numbers of the global mass matrices
Cond($M$)
m $1/h$ $\beta^{+}=10$ $\beta^{+}=100$ $\beta^{+}=1000$

### v2

@@ 4.2. Efficient generation of matrices and vectors
$\begin{split}{\color[rgb]{0,0,0}S_{K,q}}&=\sum_{s=\pm}\Big(\beta^{s}(V_{\eta}^{s})^{T}W^{s}V_{\eta}^{s}+\beta^{s}\rho^{2}(V_{\xi}^{s})^{T}W^{s}V_{\xi}^{s}\Big)\\ &=\sum_{s=\pm}\Big(\beta^{s}(C^{s})^{T}(L_{\eta}^{s})^{T}W^{s}L_{\eta}^{s}C^{s}+\beta^{s}\rho^{2}(C^{s})^{T}(L_{\xi}^{s})^{T}W^{s}L_{\xi}^{s}C^{s}\Big).\end{split}$ (4.16)
Once again, the generation of the stiffness matrix requires the evaluation of the first derivative of the polynomial basis $R_{i}$ at the quadrature points.
@@ 5. Reconstruction of Frenet-IFE Basis on Triangles
The GC-IFE bases in the general construction (4.2) focus primarily on enforcing the jump conditions, but not on computational performance when used to approximate interface problems. In this section, we introduce two reconstruction approaches that significantly improve the conditioning of the GC-IFE mass matrix and enhance computational performance.
We write Frenet IFE basis function $\hat{\lambda}_{j}$ via general construction as follows
@@
We include $C_{j}^{-},C_{j}^{+}$ in the basis functions to emphasize that these bases can be completely determined by $C_{j}^{-},C_{j}^{+}$ once the regular basis $R_{j}$ for $\mathbb{P}_{m}$ is chosen.
We aim to introduce a reconstructed Frenet IFE basis $\tilde{\lambda}$ so that the resulting mass matrix is optimally conditioned. To start, we denote the basis of $\hat{V}^{m}(\hat{K}_{F})$ by
$\hat{B}(C)=\hat{B}(C^{-},C^{+}):=\{\hat{\lambda}_{j}(\eta,\xi;C_{j}^{-},C_{j}^{+}):j=1,\cdots,d_{m}\}.$
We first recall the Lemma 1 from [2025AdjeridLinMeghaichi]:
Lemma 5.1
@@
Let $\tilde{C}=CQ$. By (4.12), the mass matrix $)M(\hat{B}(\tilde{C}$)) associated with the basis $\hat{B}(\tilde{C})=\hat{B}(\tilde{C}^{-},\tilde{C}^{+})=\hat{B}(C^{-}Q,C^{+}Q)$ can be approximated by
$\begin{split}{\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(\tilde{C}))&=\sum_{s=\pm}(\tilde{C}^{s})^{T}(L^{s})^{T}W^{s}L^{s}\tilde{C}^{s}\\ &=\sum_{s=\pm}(C^{s}Q)^{T}(L^{s})^{T}W^{s}L^{s}C^{s}Q=Q^{T}{\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))Q.\end{split}$ (5.1)
@@
We choose $Q$ such that the mass matrix becomes the identity matrix. Let $V_{1}\Lambda V_{1}^{T}={\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$ be the singular value decomposition of the matrix ${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$ where $\Lambda$ denote the diagonal matrix of positive eigenvalues of ${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$. Let $Q_{1}=V_{1}\Lambda^{-1/2}$ and $C^{(1)}=CQ_{1}$. Then we have
$\begin{split}{\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C^{(1)}))&={\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(CQ_{1}))=Q_{1}^{T}{\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))Q_{1}\\ &=\Lambda^{-1/2}V_{1}^{T}{\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))V_{1}\Lambda^{-1/2}=I_{d_{m}\times d_{m}}.\end{split}$ (5.2)
@@
The resulting approximate mass matrix ${\color[rgb]{0,0,0}M_{K,q}}(\hat{{B}}(C^{(1)}))$ achieves the optimal condition number in theory. In actual computation with finite precision, the conditioning could be affected slightly as we can see in next section.
@@
The second approach uses a factorization of ${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$ as follows. Let $n_{q}=n_{q}^{-}+n_{q}^{+}$, and combine the quadrature points and weights:
@@
then, by (4.12) we can express ${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$ as
${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))=V(\hat{B}(C))^{T}WV(\hat{B}(C)),$ (5.4)
@@
Let $\tilde{V}(\hat{B}(C))=W^{1/2}\,V(\hat{B}(C))$, then we can factorize the matrix ${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$ as follows
${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))=\tilde{V}(\hat{B}(C))^{T}\tilde{V}(\hat{B}(C)).$
@@
$\begin{split}{\color[rgb]{0,0,0}M_{K,q}}(\hat{{B}}(C^{-}Q_{2},\,C^{+}Q_{2}))&=Q_{2}^{T}{\color[rgb]{0,0,0}M_{K,q}}(\hat{{B}}(C))Q_{2}\\ &=\Sigma^{-1}V_{2}^{T}\tilde{V}^{T}\tilde{V}V_{2}\Sigma^{-1}=I_{d_{m}\times d_{m}}.\end{split}$
Both reconstruction approaches make approximate mass matrix optimally conditioned. Moreover, when the eigenvalues of ${\color[rgb]{0,0,0}M_{K,q}}(\hat{{B}}(C))$ are distinct and computations are exact, the two approaches are mathematically equivalent. In this case, they generate the same GC-IFE basis, i.e.,
@@
Remark 7.
In finite-precision computations, the two reconstruction approaches behave slightly differently. The condition number of ${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$ is the square of the condition number of $\tilde{V}(\hat{B}(C))$ since $\Lambda=\Sigma^{2}$. As a result, ${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$ is typically more ill-conditioned than $\tilde{V}$, and the numerical accuracy of the SVD in reconstruction approach (and therefore of $Q_{1}$) is limited by round-off errors in the small singular values of ${\color[rgb]{0,0,0}M_{K,q}}(\hat{B}(C))$. Consequently, the matrix $Q_{1}$ obtained numerically may deviate substantially from its exact counterpart. In contrast, the SVD of $\tilde{V}(\hat{B}(C))$ in reconstruction approach 2 is less sensitive to round-off errors because $\tilde{V}$ has a much smaller condition number. Thus the transformation matrix $Q_{2}$ is generally more accurate. These results are consistent with the reconstruction for GC-IFE basis for rectangular elements [2025AdjeridLinMeghaichi2].
@@ 6. Numerical Experiments
In this section, we present a set of numerical experiments to demonstrate the performance of the geometrically conforming immersed finite element spaces on triangular meshes.
We consider two types of computational domains. In the first three examples, the computational domain is a square and structured Cartesian triangular meshes are used. In the fourth example, we consider a hexagonal domain on which general unstructured meshes are used, see Figure 3.
In the first three examples, the computational domain is set to be a square $\Omega=(-1,1)^{2}$ with a circular interface $\Gamma=\{|\mathbf{x}|=r_{0}\}$ where $r_{0}=1/\sqrt{3}$, which partitions $\Omega$ into two subdomains $\Omega^{-}=\{|\mathbf{x}|<r_{0}\}$ and $\Omega^{+}=\{|\mathbf{x}|>r_{0}\}$. The diffusion coefficient $\beta(x)$ is piecewise constant $\beta(\mathbf{x})|_{\Omega}^{\pm}=\beta^{\pm}$ where $\beta^{-}=1$ and $\beta^{+}$ is varied in the tests.
The exact solution is given by:
$u(\mathbf{x})=\begin{cases}\dfrac{1}{\beta^{-}}\cos(\pi|\mathbf{x}|^{2}),&\mathbf{x}\in\Omega^{-},\\[5.0pt] \dfrac{1}{\beta^{+}}\cos(\pi|\mathbf{x}|^{2})+\cos(\pi r_{0}^{2})\left(\dfrac{1}{\beta^{-}}-\dfrac{1}{\beta^{+}}\right),&\mathbf{x}\in\Omega^{+},\end{cases}$ (6.1)
which satisfies the interface jump conditions (1.3)-(1.4) as well as the Laplacian extended jump condition (1.5). We examine three types of basis: $\{\hat{\lambda}_{i}\}_{i=1}^{d_{m}}$ from general construction, $\{\hat{\lambda}_{i}^{(1)}\}_{i=1}^{d_{m}}$ from the Reconstruction Approach 1, and $\{\hat{\lambda}_{i}^{(2)}\}_{i=1}^{d_{m}}$ from Reconstruction Approach 2. Polynomial degrees up to $m=9$ are tested.
@@
Figure 3. A structured triangular mesh of a rectangular domain and an unstructured mesh of a hexagonal domain.
6.1. Conditioning of Mass and Stiffness Matrices
We first investigate the conditioning of the mass and stiffness matrices associated with different choices of basis functions. Local mass matrices are computed on all interface elements of a structured $10\times 10$ triangular mesh, and the maximum condition number over all interface elements is reported in Table 1. Three types of basis functions are considered. For the general construction, the condition numbers grow rapidly as both the polynomial degree and the jump ratio increase. In contrast, both reconstruction approaches appear less sensitive to these variations. In particular, the Reconstruction Approach 2 demonstrates superior robustness at higher polynomial degrees. When the degree exceeds 7, the condition numbers associated with the Reconstruction Approach 1 increase noticeably, whereas those of the second reconstruction remain well controlled.
Table 1. Largest condition number of local mass matrices among all interface elements on a structured $10\times 10$ triangular mesh.
@@ 6.1. Conditioning of Mass and Stiffness Matrices
General RA1 RA2 General RA 1 RA 2 General RA 1 RA 2
1 5.42E+02 1 1 5.25E+04 1 1 5.25E+06 1 1
2 1.19E+05 1 1 1.24E+07 1 1 1.25E+09 1 1
3 1.47E+07 1 1 1.56E+09 1 1 1.57E+11 1 1
4 1.06E+09 1 1 1.25E+11 1 1 1.28E+13 1 1
5 4.18E+11 1 1 5.89E+13 1 1 6.58E+15 1 1
6 4.13E+13 1 1 2.67E+15 1.18 1 2.60E+17 2.59E+01 1
7 4.20E+16 1.27E+01 1 1.15E+19 1.04E+04 1 3.07E+19 2.06E+05 1
8 6.22E+18 2.22E+04 1 2.56E+19 2.02E+07 1 3.31E+21 2.91E+08 1
9 2.89E+21 1.55E+07 1 6.49E+21 5.23E+08 1 3.75E+22 1.24E+11 1
Table 2 reports the conditioning of the global mass matrices. Since the global mass matrix is block diagonal, its condition number equals the maximum condition number among all elements in the mesh. For each polynomial degree $1\leq m\leq 9$, we report the condition numbers on three different meshes. When $N=10$, i.e., using the structured $10\times 10$ triangular mesh, the condition number obtained from the general construction coincides with that in Table 1, since the largest condition number occurs on the interface elements. Moreover, no apparent growth is observed under mesh refinement. This indicates that the geometric configuration of the interface elements has a much stronger impact on the conditioning than the element size itself.
In contrast, the two reconstruction approaches yield identical condition numbers across all meshes for degrees up to $m=5$. In these cases, the condition numbers are determined entirely by non-interface elements, where an orthogonal basis [2008HesthavenWarburton] is employed. For $m\geq 6$, the condition number of the Reconstruction Approach 1 is dominated by the interface elements, whereas the Reconstruction Approach 2 remains robust. This numerical experiment confirms the Remark 7.
Table 3 reports the conditioning of the global stiffness matrices $S$ whose element components are defined in (4.14). Although the reconstruction approaches are designed to improve the conditioning of the mass matrices, they also significantly reduce the condition number of stiffness matrices.
Table 2. Condition numbers of the global mass matrix $M$ on structured $N\times N$ triangular meshes.
Cond($M$)
$m$ NN $\beta^{+}=10$ $\beta^{+}=100$ $\beta^{+}=1000$
@@
1 10 5.4E+02 4 4 5.3E+04 4 4 5.3E+06 4 4
20 4.7E+02 4 4 4.6E+04 4 4 4.6E+06 4 4
40 3.7E+02 4 4 3.6E+03 4 4 3.6E+06 4 4
2 10 1.2E+05 17 17 1.2E+07 17 17 1.3E+09 17 17
20 3.0E+05 17 17 4.0E+07 17 17 4.1E+09 17 17
40 3.5E+05 17 17 4.9E+07 17 17 5.0E+09 17 17
3 10 1.5E+07 35 35 1.6E+09 35 35 1.6E+11 35 35
20 1.1E+08 35 35 1.6E+10 35 35 1.6E+12 35 35
40 1.4E+08 35 35 2.0E+10 35 35 2.1E+12 35 35
4 10 1.0E+09 46 46 1.3E+11 46 46 1.3E+13 46 46
20 6.5E+10 46 46 9.5E+12 46 46 9.8E+14 46 46
40 8.3E+10 46 46 1.2E+13 46 46 1.3E+15 46 46
5 10 4.2E+11 62 62 5.9E+13 62 62 6.6E+15 62 62
20 1.1E+14 62 62 1.7E+16 62 62 1.8E+18 62 62
40 1.6E+14 62 62 2.3E+16 62 62 2.4E+18 62 62
6 10 2.4E+14 92 92 4.5E+16 92 92 6.4E+18 3.8E+03 92
12 2.1E+16 92 92 3.5E+18 2.2E+02 92 4.6E+20 3.2E+04 92
14 4.8E+16 92 92 6.7E+18 6.0E+02 92 8.4E+20 6.3E+04 92
7 10 4.2E+16 125 125 1.2E+19 1.0E+04 125 3.1E+19 2.1E+05 125
12 5.4E+18 1.1E+03 125 7.5E+20 3.8E+05 125 1.3E+23 2.3E+07 125
14 3.7E+19 1.0E+04 125 4.8E+21 3.7E+06 125 1.7E+24 2.5E+08 125
8 10 6.2E+18 2.2E+04 193 2.6E+19 2.0E+06 193 3.3E+21 2.9E+08 193
12 1.4E+22 1.8E+07 193 2.3E+25 1.5E+09 193 1.1E+25 1.2E+11 193
14 2.7E+22 1.3E+07 193 2.6E+24 2.2E+09 193 2.9E+26 4.7E+11 193
9 10 2.9E+21 1.6E+07 285 6.5E+21 5.2E+08 285 3.8E+22 1.2E+11 285
12 9.7E+23 2.3E+09 285 5.5E+25 3.4E+11 285 1.9E+26 2.1E+13 285
14 3.9E+24 5.3E+09 285 6.0E+26 1.4E+12 285 5.4E+29 1.0E+14 285
Table 3. Condition numbers of the global stiffness matrices on structured $N\times N$ triangular meshes.
Cond($S$)
$m$ NN $\beta^{+}=10$ $\beta^{+}=100$ $\beta^{+}=1000$